OpenStudy (anonymous):

PLEASE HELP(: Consider the infinite geometric series. a. Write the first four terms of the series. b. Does the series diverge or converge? c. If the series has a sum, find the sum.

OpenStudy (anonymous):

\[\sum_{n=1}^{infinity} -4 (1/3) ^ x-1

OpenStudy (anonymous):

\[\sum_{n=1}^{infinity}\] -4 (1/3) ^x-1

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} -4 \left(\frac{1}{3}\right)^ {x-1} \]?

OpenStudy (anonymous):

certainly converges since \(-1<\frac{1}{3}<1\)

OpenStudy (anonymous):

then use \(\frac{1}{1-\frac{1}{3}}\) to find the sum, multiply by \(-4\) at the end

OpenStudy (anonymous):

Yes that is the correct equation. Sorry I dont really know much about this website(:

OpenStudy (anonymous):

the sum would be -6, Right? And how do I use that to find the other four terms in the series?

OpenStudy (anonymous):

yes it would be -6

OpenStudy (anonymous):

the first question is a direct computation put in \(x=1\) then \(x=2\) etc

OpenStudy (anonymous):

Into the equation with the sigma in it?

OpenStudy (anonymous):

Can you tell me how to put that in the calcuator? Please(:

OpenStudy (anonymous):

no it is asking only for the terms

OpenStudy (anonymous):

\[-4\times (\frac{1}{3})^{1-1}=-4\times 0=-4\] is the first term

OpenStudy (anonymous):

rather \[-4\times (\frac{1}{3})^{1-1}=-4\times 1=-4\]

OpenStudy (anonymous):

second term, replace \(x\) by 2 and get \(-4\times \frac{1}{3}=-\frac{4}{3}\) etc

OpenStudy (anonymous):

So the third term is -4/9

OpenStudy (anonymous):

and the fourth term is -4/27? correct?

OpenStudy (anonymous):

Thank you so much(: You have been a life saver(: Do you know how trigonometric functions?

OpenStudy (anonymous):

If you think you might can please take a look at this question.(: http://openstudy.com/study#/updates/50a53189e4b0329300a8d772